OpenStudy (anonymous):

Ronald is 26 years old and starting an IRA (individual retirement account). He is going to invest $200 at the beginning of each month. The account is expected to earn 2.75% interest, compounded monthly. How much money will Ricky have in his IRA when he retires, at age 65?

OpenStudy (anonymous):

I am ASSUMING 2.75% is an ANNUAL INTEREST RATE COMPUNDED MONTHLY.. if NOT THEN PLEASE TELL ME Monthly Amount A= $200 Years n = 39 * (12 months) = 468 Rate yearly= 2.75% --- Monthly rate (r ) = 2.75/12= 0.229% = 0.229/100= 0.002292 Formula = A*(1+r) *( (1+r)n -1)/r Substituting it: = 200*(1+0.002292) *( (1+0.002292)468 -1 )/0.002292 = $ 1,67,883.68 So, after 39 years he will have this much money